OpenStudy (anonymous):

4(x-5)^2+4y^2=16 find the center(h,k) and the radius

OpenStudy (tkhunny):

Eyeball problem. Center: (5,0) -- You should see this from a quick glance. Radius is a little trickier. Divide by 4 and that should be another quick glance away.

OpenStudy (anonymous):

so I don't use (x-h)^2+(x-k)^2=r^2

OpenStudy (tkhunny):

Yes you do. Divide by 4 and it will be in that form.

OpenStudy (anonymous):

4x^2-40+4y^2=16 add 40 to both sides 4x^2+4y^2=56 divide by 4 both sides x^2+y^2=sqrt of 14 now im a little lost

OpenStudy (tkhunny):

What are you doing? What happened to that nice, convenient form. You have: \(4(x-5)^2 + 4y^{2} = 16\) Divide by 4: \((x-5)^2 + y^{2} = 4 = 2^{2}\) You are now done with the problem. Center: (5,0) Radius: 2

OpenStudy (anonymous):

I told you I was lost...guess I confused myself How do I find the intercepts?

OpenStudy (tkhunny):

That's why we invent "forms". They contain useful information. When presented with and important form, don't mess it up! Intercepts look like this: ALWAYS x-intercepts: (r,0) y-intercepts: (0,s) In words, substitute y = 0 into the equation to find x-intercepts. And, substitute x = 0 into the equation to find y-intercepts.

OpenStudy (anonymous):

is the x (2,-6)

OpenStudy (anonymous):

y=(0,14)

OpenStudy (tkhunny):

You must pay closer attention. There MUST be a zero in there, somewhere or it is NOT an intercept. Give it another go.

OpenStudy (anonymous):

when I plug in I get 14 for both x and y @tkhunny

OpenStudy (anonymous):

ok so x=(2,0) and y=(0,14)

OpenStudy (tkhunny):

You have \((x-5)^{2} + y^{2} = 4\) To find the y-intercept, substitute x = 0 \((0-5)^{2} + y^{2} = 4\) Solve for y \((5)^{2} + y^{2} = 4\) \(25 + y^{2} = 4\) \(y^{2} = 4 - 25\) \(y^{2} = -21\) Since \(y^{2}\) cannot be a negative number, there are no y-intercepts. You find the x-intercepts, if they exist.

OpenStudy (anonymous):

I added 25 and got 29

OpenStudy (tkhunny):

Why would you add 25? It's positive over there on the left. You must subtract it. Don't let your desire for a satisfying solution cloud your judgment.

OpenStudy (anonymous):

ok and I dropped my sign on the neg 5. careless mistakes

OpenStudy (anonymous):

 will my graph look like this

OpenStudy (anonymous):

or close to it

OpenStudy (tkhunny):

That's no good. The radius is WAY too big. There are NO y-intercepts. Use your x-intercepts to guide you. What are they?

OpenStudy (anonymous):

x=(2,0)

OpenStudy (anonymous):

do I point my radius first

OpenStudy (tkhunny):

You must find your x-intercepts! You have You have \((x−5)^{2} +y^{2} =4 \) To find the x-intercepts, substitute y = 0 \((x−5)^{2} +0^{2} = 4 \) \((x−5)^{2} +0 = 4 \) \((x−5)^{2} = 4 \) x - 5 = 2 or x - 5 = -2 x = 7 or x = 3 There are the x-intercepts. (7,0) and (3,0) You have to do the work. It never will be magic.